SX SOUL Checkout

2:11
  • Starting up, cooling the OCAM
  • First target is AO652
  • Preset proceeding normally
  • Hitting the Y limit on the stages, forgot to put my new pupil file in the table, I backed up the old table.
  • ncpaStartup.py works. I need to change the order in which the buttons are pressed as the ncpa trigger button doesn't like to go first (or hit it twice), but it works well otherwise
  • Star was a little too bright for full frame.
2:46
  • Going to AO658
  • 4000 frames + 10 luci frames
  • Tracking Number 20190913_025826
  • offset +10 X
  • TN 030037
  • offset +20 X
  • Off of the frame
  • Pisces dark 030508
  • -10 Y
  • For fun
  • TN: 030607
  • Back to center
  • 030747
  • I'm convinced that bin1 is working
03:10
  • AO640
  • Bin2
  • Chose the wrong guide star and the re-rotator hit its limit,
  • Homed the rotator just for good measure.
  • Had to chose PA of 90, GS=8
  • Going to try GOPT and NCPA
  • TN 033126, 033237
  • Offset of +25 x, to get it off of the field
  • Pisces dark: 033547
  • Now I'm going to do a PA of 270, to see if the pupils can be centered by the camera lens
  • Need guide star number 19
  • It centered:
  • one TN: 034349
  • Bin2 works with NCPA, at least in subarcsecond seeing
03:45
  • Going to try for a faint, bin 4 star
  • AO664
  • The pupil position could stand to be tweaked, We are at +100X, but the pupil is 0.31 off of being centered.
  • in the interest of time, I'll rotate the instrument by 180, pa=270
  • It centers here
  • I'll lengthen the LUCI exposure to 20 seconds
  • As I recall, Alfio was having trouble with optimizing the gain at this step, but we'll try
  • TN: 040052 Error! It gives an IO error, every few seconds, 10, 20 second frames made it go over the limit of 40 tries
  • TN: only 4 frames: 040509
  • Judging by the LUCI frame, I'll offset by -25 y, to clear the three stars that are in the frame.
  • Pisces dark: 041348
  • TN: 041422
04:16
  • I'm going to test NCPA a bit, in bin 1 and look at the camera lens centering
  • AO735
  • Starting PA=0
  • TN: 042524
  • Offset y=25
  • Pisces dark 042839
  • rot offset of +30
  • 043006
  • rot offset of +60 (87 degrees on LUCI rotator)
  • 043202
  • rot offset of +90 (115 on luci, going down)
  • 043504
  • Changing guide star to get +120
  • 044406
  • +150
04:47
  • Adsec went into failure, unknown cause
  • +150
  • 045441
  • +180
  • 045706
  • +210
  • 045905
  • +240
  • Approaching the X limit!!! 96.24
  • 050131
  • +270
  • Went back to 83.2
  • 050343
  • +300
  • 051054
  • +330
  • 051304
  • Since the instrument rotator is moving toward the negative, I'll go a few more degrees, to make sure that I cover the whole 360 degrees
  • +360
  • 052057
  • +390
  • 052255
  • Pisces dark
-- GregoryTaylor - 13 Sep 2019
Topic revision: r2 - 13 Sep 2019, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback